OpenStudy (isaac.tillman03):

Solve for X and express in simplest form 1/7 = 3/ 4x -1

OpenStudy (isaac.tillman03):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

well you wanna get x by itself therefore add one to both sides so 1/7+1=3/4x-1+1 which cancels those out so 1 1/7=3/4x

rebeccaxhawaii (rebeccaxhawaii):

make it a improper fraction 8/7=3/4x

OpenStudy (isaac.tillman03):

so is the answer 21/32 cause i got that but i wasn't sure.

rebeccaxhawaii (rebeccaxhawaii):